a $45 shirt is on sale for 1/5 off. How much is the sales price?
Found 2 solutions by nyc_function, Alan3354:
Answer by nyc_function(2741)

Let s = sales price
1/5 = 0.20
s = 45 * 0.20
s = $9


Answer by Alan3354(69443)

1/5 off leaves 4/5 of the $45.
(4/5)*$45 = $36